# Obfuscation-Resilient Binary Code Similarity Analysis using Dominance Enhanced Semantic Graph

### v1

@@ Abstract.
Binary code similarity analysis (BCSA) serves as a core technique for binary analysis tasks such as vulnerability detection. While current graph-based BCSA approaches capture substantial semantics and show strong performance, their performance suffers under code obfuscation due to the unstable control flow. To address this issue, we develop ORCAS, an Obfuscation-Resilient BCSA model based on Dominance Enhanced Semantic Graph (DESG). The DESG is an original binary code representation, capturing more binaries’ implicit semantics without control flow structure, including inter-instruction relations, inter-basic block relations, and instruction-basic block relations. ORCAS robustly scores semantic similarity across binary functions from different obfuscation options, optimization levels, and instruction set architectures. Extensive evaluation on the BinKit dataset shows ORCAS significantly outperforms eight baselines, achieving an average 12.1% PR-AUC gain when using combined three obfuscation options compared to the state-of-the-art approaches. Furthermore, ORCAS improves recall by up to 43% on an original obfuscated real-world vulnerability dataset, which we released to facilitate future research.
Binary Code Similarity Analysis, Obfuscation-Resilient, Dominator Tree
ccs: Security and privacy Software reverse engineeringccs: Computing methodologies Machine learning
@@ 1. Introduction
Code obfuscation obscures the code’s control flow and basic blocks while preserving the function semantics, which has been widely applied to increase the difficulty of unauthorized reverse engineering (Gao et al., 2023), e.g., in 2018, roughly 50% of popular apps on Google Play with over 10 million downloads were obfuscated to protect against plagiarism and repackaging (Wermke et al., 2018). While malware authors also exploit code obfuscation techniques to conceal malicious contents and evade detection (Cheng et al., 2021; Cuiying et al., 2024). In 2024, approximately eight million new malware samples are identified, with over 300,000 being detected each day. [^1] https://portal.av-atlas.org/malware
@@
Binary code similarity analysis (BCSA) scores the semantic similarity between two given binary code snippets, which is commonly used for malware identification (Cesare et al., 2013; Kim et al., 2019), vulnerability detection (Luo et al., 2023; Li et al., 2023), patch analysis (Hu et al., 2016; Xu et al., 2017a; Kargén and Shahmehri, 2017), and software plagiarism detection (Luo et al., 2017), etc. Existing BCSA approaches perform well in BCSA tasks across different compilers, optimization levels, and instruction set architectures (ISAs). However, they face challenges when analyzing obfuscated binary code. To be specific, a binary function can be represented as a control flow graph (CFG), where a node represents a basic block, and the directed edge between two nodes represents their control flow. There are three widely used code obfuscation techniques: instruction substitution (SUB), bogus control flow (BCF), and control flow flattening (FLA) (Junod et al., 2015). SUB obfuscation replaces simple instructions with semantically equivalent complex instructions, with control flow untouched. BCF obfuscation modifies the control flow by adding bogus basic blocks and edges. While FLA obfuscation complicates the control flow with a complex hierarchy, and then flattens the control flow into a linear form. Meanwhile, different techniques can be combined to obfuscate the code, e.g., for a given CFG, first, BCF is applied to double the number of nodes and edges, then FLA is applied to double them again (Ding et al., 2019). Then we can see that a CFG and its obfuscated CFG can have different nodes, different flow structures, and even arbitrarily added bogus nodes and edges, which present great challenges to existing BCSA approaches to detect their semantic similarity.
@@
According to how the binary code is represented, existing BCSA approaches can be categorized into instruction stream based and graph based. When instruction stream based approaches are used, binary code is first translated into assembly code or intermediate representation (IR) code, which are treated as natural language sentences. Then natural language processing (NLP) technologies are applied to capture the code’s implicit semantics. NLP approaches have strong contextual understanding capabilities, however, they usually rely on pre-training tasks tailored to binary code, which require large-scale datasets and expensive computational resources (Wang et al., 2022; Li et al., 2021; Ahn et al., 2022; Yu et al., 2020; Luo et al., 2023). In addition, obfuscating simple instructions with complex ones increases contextual complexity and token length, which degrades the performance of NLP technologies.
@@
When graph based approaches are applied, a binary function is represented as a feature enhanced CFG or semantics-oriented graph (SOG). Then graph neural network (GNN) technologies are used to learn the code’s semantics. First, various features are incorporated into the CFG for capturing code’s semantics: (1) Incorporating basic block features (e.g., number of instructions), which are obtained by manually crafting (Xu et al., 2017b; Feng et al., 2016) or automatically generating using NLP technologies (Luo et al., 2023; Yu et al., 2020; Jia et al., 2024a); (2) Incorporating data flow graph (DFG) and call graph (CG) to provide richer semantic information between instructions and functions (Zhang et al., 2024; Duan et al., 2020; Guo et al., 2022). Second, when a binary function is represented as a SOG (He et al., 2024), nodes represent instruction opcodes and operands and edges reveal intra-instruction as well as inter-instruction relations, where control flow between basic blocks are represented using inter-instruction relations. Graph based approaches capture more binary code semantics and demonstrate remarkable BCSA performance. However, when code obfuscation technologies are applied, the unstable control flow structure degrades graph based BCSA performance (Ding et al., 2019), this problem remains unresolved.
@@
The rest of the paper is organized as follows. Section 2 introduces our motivation. Section 3 presents our original BCSA model ORCAS. Section 4 describes the experimental setup. Section 5 describes the performance evaluation of ORCAS. Section 6 reviews the related work and finally Section 7 concludes the paper.
@@ 2. Motivation
This section describes the motivation behind this paper. In general, BCSA can be conducted across different granularity, ranging from coarse to fine: program-level, function-level, slice-level, basic block-level, and instruction-level (Haq and Caballero, 2021). This paper focus on the function-level BCSA since functions are considered basic building blocks of program functionality, function-level BCSAs (Wang et al., 2022; Xu et al., 2023; Luo et al., 2023; He et al., 2024; Zou et al., 2022; Ding et al., 2019; Pei et al., 2023; Feng et al., 2024) enables many real-world applications such as patch analysis (Feng et al., 2019; Yang et al., 2023b), and can be easily extended to detect similarities between whole programs.
Definition 2.1 (Function).
A function is an ordered sequence of binary instructions compiled from a function written in high-level programming languages, based on different settings, e.g., obfuscation options, optimization levels, and ISAs. A binary function possesses specific semantics and structures, which can be represented as assembly code, IR code, a graph (e.g., CFG), or a tree (e.g., abstract syntax tree).
Actually, a binary function can be represented as a CFG, denoted as $G=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}=\{v_{0},v_{1},\ldots,v_{n-1}\}$ is the set of nodes, $v_{0}$ and $v_{n-1}$ denotes the unique entry and exit node respectively. =$\mathcal{E}=\{(v_{i},v_{j})\mid v_{i},v_{j}\in\mathcal{V}\}$ is the set of directed edges between nodes. To be specific, a node represents a basic block, which is a sequence of consecutive instructions ending with a branch or return instruction.
To put the discussion into perspective, an example CFG $G$ is depicted in Fig. LABEL:fig:motivation(a) to illustrate the problem. SUB obfuscates instructions, i.e., the nodes in a CFG, while FLA and BCF obfuscate nodes as well as edges in a CFG, which are widely adopted by various BCSA approaches (He et al., 2024; Li et al., 2021; Guo et al., 2022; David et al., 2017; Jia et al., 2024b; Luo et al., 2023) for more notable performances. Then, BCF and FLA obfuscation have been applied to CFG $G$ and obtained an obfuscated CFG $\tilde{G}$, which is depicted in Fig. LABEL:fig:motivation(d), where the dashed boxes display a typical application of an official Obfuscator-LLVM’s BCF and FLA obfuscation techniques (Junod et al., 2015). We can see that node $v_{3}$ and node $v_{4}$ have an edge to node $v_{5}$ in CFG $G$, while in the obfuscated CFG $\tilde{G}$, neither node $v_{3}$ nor node $v_{4}$ has an edge to node $v_{5}$. That is, the control flow structure of CFG $G$ is different from that of CFG $\tilde{G}$.
@@
Based on the aforementioned discussion, we observe that dominance and post-dominance relations exhibit superior structural stability compared to control flow structure under obfuscation. To quantify this stability difference, we employ graph edit distance (GED), similar to (Pawlowski et al., 2016), which measures the minimum number of edit operations (i.e., node/edge additions/deletions or node relabelling) required to transform one graph into another. Specifically, we compute GED between original and obfuscated versions of both CFGs and D-Trees using approximately 1,800 function samples obfuscated with combined BCF and FLA obfuscation. Table 1 presents a comparative analysis of structural stability between CFG and D-Tree under obfuscation across sizes of basic blocks. Notably, D-Trees exhibit a 23.6% lower average GED for functions with basic block sizes exceeding 150. This demonstrates that D-Tree consistently exhibit superior structural stability compared to CFG under obfuscation.
@@ 3. ORCAS
This section elaborates on how ORCAS is designed. Fig. 2 depicts the ORCAS model, which takes two binary functions as input and decompiles them using Ghidra (Eagle and Nance, 2020) into P-Code IR functions. In the function embedding phase, we construct the DESGs and employ a Gated Graph Neural Network (GGNN) (Li et al., 2015) to generate embeddings. In the similarity computation phase, we utilize cosine similarity to measure binary function pairs’ similarity using their embeddings.
@@ 3.1. Intermediate Representation
To mitigate the differences between various ISAs and simplify instruction types, IRs are widely utilized to represent binary functions in recent BCSA approaches (Yu et al., 2023; He et al., 2024; Luo et al., 2023). Since Ghidra outperforms other decompilers like IDA-Pro (Eagle, 2011) and Angr (Shoshitaishvili et al., 2016) in correctly identified function starts task (Shaila et al., 2021), we utilize Ghidra to decompile the binary code and lift it into P-Code, Ghidra’s IR.
@@
Table 2. Varnode normalization strategies.
@@
  • For ram address space, if the varnode represents a standard library function, we normalize it using the function name; if not, we normalize it with the identifier ram. Notably, standard library function retain their original function name by default under obfuscation, as symbol table-based dynamic linking would fail with obfuscated name (Picheta, 2020).
@@ 3.2.1. Implicit Semantics of Binary Function
Inter-basic block relations. Control flow relations define the original relations between basic blocks in a CFG. However, bogus control flow generated by obfuscation techniques may introduce misleading information. As a result, we chose to remove them. Based on CFG, we introduce two new relations between basic blocks: dominance and post-dominance relations, derived from dominance analysis (Lengauer and Tarjan, 1979; Azen and Budescu, 2003). These relations are crucial for understanding control flow structure and analyzing the execution paths of obfuscated binary function.
Inter-instruction relations. We further divide inter-instruction relations into two categories: data relation and effect relation. In previous binary analysis works (Li et al., 2021; David et al., 2017; He et al., 2024; Guo et al., 2022; Jia et al., 2024b; Luo et al., 2023), the def-use relation between instructions is often used since it is a fundamental concept in program analysis. We adopt the def-use relation as the data relation, which reveals how instruction use values are defined by others. Inspired by (Click and Paleczny, 1995; He et al., 2024), effect relation imposes restrictions on the execution order between instructions, which is derived from def-use relations and it is essentially a special case of the def-use relation. Especially the effect relation focuses on the execution order between a CALL instruction and an instruction that may read from or write to memory. Take Fig. 3a as an example, the instruction r2=abs(r1) may use the value r1 defined by the memory read instruction r1=LOAD(0x8). This establishes an effect relation between them, indicating that it must be ensured that the instruction r1=LOAD(0x8) executes before the instruction r2=abs(r1).
Instruction-basic block relations. A CFG inherently implies that instructions are contained within specific basic blocks. To explicitly expose this to the GNN, we introduce a contain relation between basic blocks and instructions, indicating that an instruction belongs to a particular basic block. This relation directly reduces the distance between instructions within the same basic block in the graph, facilitating information interaction during the GNN training phase.
@@ 3.2.2. Steps to the DESG
Step ❸. Based on the ISG, we further reveal the structure within all instructions by splitting the opcode and operand(s) (i.e., varnodes) into tokens. As a result, the relations between instructions are refined into relations between tokens, with the effect relation specifically defined on the opcode tokens. Similar to (He et al., 2024), the relation between opcode tokens and operand tokens can be considered a data relation, which we refine into two distinct relations: one for the first operand and one for the second operand (if exists). For example, in the instruction x86_r_0 = INT_SUB x86_r_0, c_0, the tokens x86_r_0 and c_0 are the first and the second operand of the token INT_SUB, respectively. Correspondingly, there are data relations between them. Consequently, we obtain the token-based semantic graph (TSG).
@@ 3 Apply varnode normalization strategies on $\mathcal{G}_{\text{CFG}}$;
7             $\mathcal{G}_{\text{ISG}}$.addNode($inst$);
@@
10                   if $op$ is defined by instruction $defInst$ then
11                         $\mathcal{G}_{\text{ISG}}$.addDataEdge($inst$, $defInst$);
12                         if $defInst$ accesses memory then
13                               $\mathcal{G}_{\text{ISG}}$.addEffectEdge($inst$, $defInst$);
@@
17       Split the opcode of $inst$ as a node $opc$;
18       Refine the effect edge of $inst$ into the effect edge of $opc$;
19       Split the target operand of $inst$ as a node $op$;
20       $\mathcal{G}_{\text{TSG}}$.addDataEdge($op$, $opc$);

### v2

# ORCAS: Obfuscation-Resilient Binary Code Similarity Analysis using Dominance Enhanced Semantic Graph

@@ Abstract.
Binary code similarity analysis (BCSA) serves as a foundational technique for binary analysis tasks such as vulnerability detection and malware identification. Existing graph based BCSA approaches capture more binary code semantics and demonstrate remarkable performance. However, when code obfuscation is applied, the unstable control flow structure degrades their performance. To address this issue, we develop ORCAS, an Obfuscation-Resilient BCSA model based on Dominance Enhanced Semantic Graph (DESG). The DESG is an original binary code representation, capturing more binaries’ implicit semantics without control flow structure, including inter-instruction relations (e.g., def-use), inter-basic block relations (i.e., dominance and post-dominance), and instruction-basic block relations. ORCAS takes binary functions from different obfuscation options, optimization levels, and instruction set architectures as input and scores their semantic similarity more robustly. Extensive experiments have been conducted on ORCAS against eight baseline approaches over the BinKit dataset. For example, ORCAS achieves an average 12.1% PR-AUC improvement when using combined three obfuscation options compared to the state-of-the-art approaches. In addition, an original obfuscated real-world vulnerability dataset has been constructed and released to facilitate a more comprehensive research on obfuscated binary code analysis. ORCAS outperforms the state-of-the-art approaches over this newly released real-world vulnerability dataset by up to a recall improvement of 43%.
Binary Code Similarity Analysis; Obfuscation-Resilient; Dominator Tree
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 34th ACM International Conference on Information and Knowledge Management; November 10–14, 2025; Seoul, Republic of Koreabooktitle: Proceedings of the 34th ACM International Conference on Information and Knowledge Management (CIKM ’25), November 10–14, 2025, Seoul, Republic of Koreadoi: 10.1145/3746252.3761266isbn: 979-8-4007-2040-6/2025/11ccs: Security and privacy Software reverse engineeringccs: Computing methodologies Machine learning
@@ 1. Introduction
Code obfuscation obscures the code’s control flow and basic blocks while preserving the function semantics, which has been widely applied to increase the difficulty of unauthorized reverse engineering (gao2023obfuscation), e.g., in 2018, roughly 50% of popular apps on Google Play with over 10 million downloads were obfuscated to protect against plagiarism and repackaging (wermke2018). While malware authors also exploit code obfuscation techniques to conceal malicious contents and evade detection (obfuscation2021; cuiying2024). In 2024, approximately eight million new malware samples are identified, with over 300,000 being detected each day. [^1] https://portal.av-atlas.org/malware
@@
Binary code similarity analysis (BCSA) scores the semantic similarity between two given binary code snippets, which is commonly used for malware identification (control2013; kim2019), vulnerability detection (vulhawk2023; sense2023), patch analysis (cross2016; spain2017; kargen2017), and software plagiarism detection (semantics2017), etc. Existing BCSA approaches perform well in BCSA tasks across different compilers, optimization levels, and instruction set architectures (ISAs). However, they face challenges when analyzing obfuscated binary code. To be specific, a binary function can be represented as a control flow graph (CFG), where a node represents a basic block, and the directed edge between two nodes represents their control flow. There are three widely used code obfuscation techniques: instruction substitution (SUB), bogus control flow (BCF), and control flow flattening (FLA) (obfuscator2015). SUB obfuscation replaces simple instructions with semantically equivalent complex instructions, with control flow untouched. BCF obfuscation modifies the control flow by adding bogus basic blocks and edges. While FLA obfuscation complicates the control flow with a complex hierarchy, and then flattens the control flow into a linear form. Meanwhile, different techniques can be combined to obfuscate the code, e.g., for a given CFG, first, BCF is applied to double the number of nodes and edges, then FLA is applied to double them again (asm2vec2019). Then we can see that a CFG and its obfuscated CFG can have different nodes, different flow structures, and even arbitrarily added bogus nodes and edges, which present great challenges to existing BCSA approaches to detect their semantic similarity.
@@
According to how the binary code is represented, existing BCSA approaches can be categorized into instruction stream based and graph based. When instruction stream based approaches are used, binary code is first translated into assembly code or intermediate representation (IR) code, which are treated as natural language sentences. Then natural language processing (NLP) technologies are applied to capture the code’s implicit semantics. NLP approaches have strong contextual understanding capabilities, however, they usually rely on pre-training tasks tailored to binary code, which require large-scale datasets and expensive computational resources (jtrans2022; palmtree2021; binshot2022; order2020; vulhawk2023). In addition, obfuscating simple instructions with complex ones increases contextual complexity and token length, which degrades the performance of NLP technologies.
@@
When graph based approaches are applied, a binary function is represented as a feature enhanced CFG or semantics-oriented graph (SOG). Then graph neural network (GNN) technologies are used to learn the code’s semantics. First, various features are incorporated into the CFG for capturing code’s semantics: (1) Incorporating basic block features (e.g., number of instructions), which are obtained by manually crafting (gemini2017; genius2016) or automatically generating using NLP technologies (vulhawk2023; order2020; jia2024cross); (2) Incorporating data flow graph (DFG) and call graph (CG) to provide richer semantic information between instructions and functions (semantics2024; deepbindiff2020; guo2022). Second, when a binary function is represented as a SOG (sog2024), nodes represent instruction opcodes and operands and edges reveal intra-instruction as well as inter-instruction relations, where control flow between basic blocks are represented using inter-instruction relations. Graph based approaches capture more binary code semantics and demonstrate remarkable BCSA performance. However, when code obfuscation technologies are applied, the unstable control flow structure degrades graph based BCSA performance (asm2vec2019), this problem remains unresolved.
@@ 2. Motivation
This section describes the motivation behind this paper. In general, BCSA can be conducted across different granularity, ranging from coarse to fine: program-level, function-level, slice-level, basic block-level, and instruction-level (ASurveyofBinaryCodeSimilarity2021). This paper focus on the function-level BCSA since functions are considered basic building blocks of program functionality, function-level BCSAs (jtrans2022; xu2023improving; vulhawk2023; sog2024; mvulpreter2022; asm2vec2019; Trex2023; crabs-former2024) enables many real-world applications such as patch analysis (PATCHDETECTOR2019; Robin2023), and can be easily extended to detect similarities between whole programs.
Actually, a binary function can be represented as a CFG, denoted as $G=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}=\{v_{0},v_{1},\ldots,v_{n-1}\}$ is the set of nodes, $v_{0}$ and $v_{n-1}$ denote the unique entry and exit node respectively. $\mathcal{E}=\{(v_{i},v_{j})\mid v_{i},v_{j}\in\mathcal{V}\}$ is the set of directed edges between nodes. To be specific, a node represents a basic block, which is a sequence of consecutive instructions ending with a branch or return instruction.
To put the discussion into perspective, an example CFG $G$ is depicted in Fig. LABEL:fig:motivation(a) to illustrate the problem. SUB obfuscates instructions, i.e., the nodes in a CFG, while FLA and BCF obfuscate nodes as well as edges in a CFG, which are widely adopted by various BCSA approaches (sog2024; palmtree2021; guo2022; similarity2017; codeextract2024; vulhawk2023) for more notable performances. Then, BCF and FLA obfuscation have been applied to CFG $G$ and obtained an obfuscated CFG $\tilde{G}$, which is depicted in Fig. LABEL:fig:motivation(d), where the dashed boxes display a typical application of an official Obfuscator-LLVM’s BCF and FLA obfuscation techniques (obfuscator2015). We can see that node $v_{3}$ and node $v_{4}$ have an edge to node $v_{5}$ in CFG $G$, while in the obfuscated CFG $\tilde{G}$, neither node $v_{3}$ nor node $v_{4}$ has an edge to node $v_{5}$. That is, the control flow structure of CFG $G$ is different from that of CFG $\tilde{G}$.
@@
Based on the aforementioned discussion, we observe that dominance and post-dominance relations exhibit superior structural stability compared to control flow structure under obfuscation. To quantify this stability difference, we employ graph edit distance (GED), similar to (ged2016), which measures the minimum number of edit operations (i.e., node/edge additions/deletions or node relabelling) required to transform one graph into another. Specifically, we compute GED between original and obfuscated versions of both CFGs and D-Trees using approximately 1,800 function samples obfuscated with combined BCF and FLA obfuscation. Table 1 presents a comparative analysis of structural stability between CFG and D-Tree under obfuscation across sizes of basic blocks. Notably, D-Trees exhibit a 23.6% lower average GED for functions with basic block sizes exceeding 150. This demonstrates that D-Tree consistently exhibit superior structural stability compared to CFG under obfuscation.
@@ 3. ORCAS
This section elaborates on how ORCAS is designed. Fig. 2 depicts the ORCAS model, which takes two binary functions as input and decompiles them using Ghidra (ghidra2020) into P-Code IR functions. In the function embedding phase, we construct the DESGs and employ a Gated Graph Neural Network (GGNN) (GGNN2015) to generate embeddings. In the similarity computation phase, we utilize cosine similarity to measure binary function pairs’ similarity using their embeddings.
@@ 3.1. Intermediate Representation
To mitigate the differences between various ISAs and simplify instruction types, IRs are widely utilized to represent binary functions in recent BCSA approaches (cfg2vec2023; sog2024; vulhawk2023). Since Ghidra outperforms other decompilers like IDA-Pro (ida2011) and Angr (angr2016) in correctly identified function starts task (disco2021), we utilize Ghidra to decompile the binary code and lift it into P-Code, Ghidra’s IR.
@@
Table 2. An example of varnode normalization strategies.
@@
  • For ram address space, if the varnode represents a standard library function, we normalize it using the function name; if not, we normalize it with the identifier ram. Notably, standard library function retain their original function name by default under obfuscation, as symbol table-based dynamic linking would fail with obfuscated name (picheta2020).
@@ 3.2.1. Implicit Semantics of Binary Function
Inter-basic block relations. Control flow relations define the original relations between basic blocks in a CFG. However, bogus control flow generated by obfuscation techniques may introduce misleading information. As a result, we chose to remove them. Based on CFG, we introduce two new relations between basic blocks: dominance and post-dominance relations, derived from dominance analysis (lengauer1979fast; azen2003dominance). These relations are crucial for understanding control flow structure and analyzing the execution paths of obfuscated binary function.
Inter-instruction relations. We further divide inter-instruction relations into two categories: data relation and effect relation. In previous binary analysis works (palmtree2021; similarity2017; sog2024; guo2022; codeextract2024; vulhawk2023), the def-use relation between instructions is often used since it is a fundamental concept in program analysis. We adopt the def-use relation as the data relation, which reveals how instruction use values are defined by others. Inspired by (click1995simple; sog2024), effect relation imposes restrictions on the execution order between instructions, which is derived from def-use relations and it is essentially a special case of the def-use relation. Especially the effect relation focuses on the execution order between a CALL instruction and an instruction that may read from or write to memory.
Instruction-basic block relations. A CFG inherently implies that instructions are contained within specific basic blocks. To explicitly expose this to the GNN, we introduce a contain relation between basic blocks and instructions, indicating that an instruction belongs to a particular basic block. This relation directly reduces the distance between instructions within the same basic block in the graph, facilitating information interaction within the GNN.
@@ 3.2.2. Steps to the DESG
Step ❸. Based on the ISG, we further reveal the structure within all instructions by splitting the opcode and operand(s) (i.e., varnodes) into tokens. As a result, the relations between instructions are refined into relations between tokens, with the effect relation specifically defined on the opcode tokens. Similar to (sog2024), the relation between opcode tokens and operand tokens can be considered a data relation, which we refine into two distinct relations: one for the first operand and one for the second operand (if exists). Consequently, we obtain the token-based semantic graph (TSG).
@@ 3 Apply varnode normalization strategies on $\mathcal{G}_{\text{CFG}}$;
7     $\mathcal{G}_{\text{ISG}}$.addNode(instinst);
@@
10        if opop is defined by instruction defInstdefInst then
11           $\mathcal{G}_{\text{ISG}}$.addDataEdge(instinst, defInstdefInst);
12           if defInstdefInst accesses memory then
13              $\mathcal{G}_{\text{ISG}}$.addEffectEdge(instinst, defInstdefInst);
@@
17  Split the opcode of instinst as a node opcopc;
18  Refine the effect edge of instinst into the effect edge of opcopc;
19  Split the target operand of instinst as a node opop;
20  $\mathcal{G}_{\text{TSG}}$.addDataEdge(opop, opcopc);